MY SCHOOL IS 12.531 KM FROM MY HOUSE. wHAT DISTANCE WILL I TRAVEL IF I GO AND COME BACK 8 TIMES
Answer by geetha_rama(94)

Ditance from house to school = 12.531 KM
Distance from house to school and back to home = 2*12.531 = 25.062 KM
total Distance for 8 times (going to school and back home) = 25.062 * 8
= 200.496 KM